what is the solution to 4/9=r-3/6
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

4/9=(r-3)/6

4*6=9(r-3)

24=9r-27

24+27=9r

51=9r

51/9=r

17/3=r

r=17/3

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

Answer by Alan3354(69443)

You should say "solve for r"
4/9=r-3/6
r = 4/9 + 1/2
r = 8/18 + 9/18
r = 17/18